A new operating system for an existing machine is expected to cost $740,000 and have a...

  1. A new operating system for an existing machine is expected to cost $740,000 and have a useful life of six years. The system yields an incremental after-tax income of $215,000 each year after deducting its straight-line depreciation. The predicted salvage value of the system is $25,000.
  2. A machine costs $380,000, has a $33,500 salvage value, is expected to last eight years, and will generate an after-tax income of $84,000 per year after straight-line depreciation.

Assume the company requires a 12% rate of return on its investments. Compute the net present value of each potential investment. (PV of $1, FV of $1, PVA of $1, and FVA of $1) (Use appropriate factor(s) from the tables provided.)

A new operating system for an existing machine is expected to cost $740,000 and have a useful life of six years. The system yields an incremental after-tax income of $215,000 each year after deducting its straight-line depreciation. The predicted salvage value of the system is $25,000. (Round your answers to the nearest whole dollar.)

Cash Flow Select Chart Amount x PV Factor = Present Value
Annual cash flow = $0
Residual value = 0
Net present value

A machine costs $380,000, has a $33,500 salvage value, is expected to last eight years, and will generate an after-tax income of $84,000 per year after straight-line depreciation. (Round your answers to the nearest whole dollar.)

Cash Flow Select Chart Amount x PV Factor = Present Value
Annual cash flow = $0
Residual value = 0
Net present value
Answer #1

L

a) Annual Depreciation = ($740000-25000)/6 years =$119167
Annual Cash Flow = Net Income + Annual Depreciation
=$215000+119167
=$334167
Cash Flow Select Chart Amount * PV Factor = Present Value
Annual Cash Flow PVA of $1 $                                              3,34,167 * 4.111407 = $         13,73,897
Residual Value PV of $1 $                                                  25,000 * 0.506631 = $               12,666
Initial Investment $            7,40,000
Net Present Value $            6,46,562
a) Annual Depreciation = ($380000-33500)/8 years =$43312.5
Annual Cash Flow = Net Income + Annual Depreciation
=$84000+43312.50
=$127312.5
Cash Flow Select Chart Amount * PV Factor = Present Value
Annual Cash Flow PVA of $1 $                                              1,27,313 * 4.96764 = $            6,32,443
Residual Value PV of $1 $                                                  33,500 * 0.403883 = $               13,530
Initial Investment $      3,80,000.00
Net Present Value $            2,65,973

Let me know wrong answer,if any.